alberto wants to make 10 liters of a 13% solution by mixing together a 10% acid solution and a 20% acid solution. how much of each solution must he use ?
Answer by TimothyLamb(4379)

---
10x + 20y = 13*10
x + y = 10
---
put the system of linear equations into standard form
---
10x + 20y = 130
x + y = 10
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 10% solution 7 liters
y= 20% solution 3 liters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php